Suppose that f(x)=9x4/5−2x5/7. Evaluate each of the following:

f′(3) =

f′(6) =
1
Expert's answer
2020-11-09T20:25:13-0500

Given "f(x)=9x^{\frac{4}{5}}-2x^{\frac{5}{7}}" then;


"f'(x)=\frac{36}{5}x^{-\frac{1}{5}}-\frac{10}{7}x^{-\frac{2}{7}}=\frac{36}{5x^{\frac{1}{5}}}-\frac{10}{7x^{\frac{2}{7}}}"


"f'(x)=\frac{36}{5x^{\frac{1}{5}}}-\frac{10}{7x^{\frac{2}{7}}}" .


Then


"f'(3)=\frac{36}{5(3)^{\frac{1}{5}}}-\frac{10}{7(3)^{\frac{2}{7}}} \approx4.736025022"


"f'(6)=\frac{36}{5(6)^{\frac{1}{5}}}-\frac{10}{7(6)^{\frac{2}{7}}} \approx 4.175359577"